12 GeV Upgrade
The performance of the C100 cryomodule has now been successfully demonstrated at 104 Megavolts for more than one hour. Preparations are complete for a test at 108 MV with full beam loading this week. Assembly of the sixth C100 has been completed. Installation of the compressors and 4.5 Kelvin coldbox for the Central Helium Liquefier-2 continues. The beam position monitors for arc 10 are ready for installation. The first-article electronics for the stripline beam position monitors passed acceptance testing.
Hall A ran beautifully all week, collecting a lot of data. One more week to go before the shutdown.
In addition to data taking on liquid hydrogen (LH2) and aluminum targets, the Q-weak experiment in Hall C has made several background measurements and carried out several test plans that will aid in analysis of the data. Although Hall C experienced significant flooding from Tuesday's storm, there was no apparent equipment damage, and Hall C is expected to receive beam for the last two days of 6 GeV operations.
The week began with a high rate of beam interruptions due to accelerating cavities tripping off on various faults. Several cavities in the North Linac were restored to service in an effort to lower the trip rate. Wednesday was a beam studies day, with tests for the Polarized Electrons for Polarized Positrons experiment and the C100 cryomodules being performed. Some more cavities in the North Linac were also recovered and brought online again, helping to lower the beam trip rate. Thursday saw a small vacuum problem corrected by an adjustment of the setpoint at which the vacuum valve closes. On Friday, a controller that handles magnets was misidentified and was inadvertently rebooted, which resulted in the loss of a few hours to restore the magnets and correct the resulting orbit and beam optics. The final weekend this year of running the main accelerator went extremely well, with no major interruptions to the beam.
First lasing on the visible ultraviolet system was achieved with the new wiggler. Substantial gain was observed. The system is being optimized and minor hardware failures are being fixed in preparation for user experiments later this month. The FEL team also began hardware installation preparations for the DarkLight effort planned for July.
The Machine Shop is presently manufacturing or continuing to manufacture:
The shop also received the new vertical milling machining center that is needed for reworking magnet cores associated with the 12 GeV Upgrade. The shop is in the process of setting the machine up. The shop has processed 156 job requests since April 2.
Environment, Safety, Health and Quality
Verify or Die: Never Assume Circuits Are Off!
If all of the steps are performed correctly, verification becomes a non-event. If, however, one error is made in any of the prescribed steps, verification of zero energy can save your life. Visit the link for a complete toolbox talk on this topic.
Computing and Networking Infrastructure
Email Migrations into New System Continue
CNI Maintenance Period Set for May 22
JLab 2012 T-Shirts Available for Purchase
DOE Lab Breakthroughs in 140 Characters or Less
Tracking Thomas at JLab
Congratulations this week go to Russell Mammei, who was the first to correctly identify the May 9 location. An honorable mention goes to Harry Fanning. Check out the Tracking Thomas webpage for a better view of his last location and this week's new mystery photo.
JLab Calendar of Events
May 14-17: QCD Evolution Workshop